OpenStudy (tayfishb):

find a point c satisfying the conclusion of the mean value theorem for the given function and interval y=x/(x+2) [1,4]

OpenStudy (tayfishb):

\[f \prime x=\frac{ -x }{ (x+2)^{2} }\]

OpenStudy (tayfishb):

\[\frac{ \frac{ 4 }{ 4+2 }-\frac{ 1 }{ 1+2 } }{4-1 } \]

OpenStudy (tayfishb):

\[\frac{ \frac{ 2 }{ 3 } -\frac{ 1 }{3 }}{3 }\]

OpenStudy (tayfishb):

\[\frac{ 1 }{ 3 }*\frac{ 1 }{ 3 }=\frac{ 1 }{ 9 }\]

OpenStudy (tayfishb):

\[\frac{ -x }{ (x+2)^{2} }=\frac{ 1 }{9 }\]

OpenStudy (tayfishb):

\[-9x=x ^{2}+4x+4\]

OpenStudy (tayfishb):

\[0=x ^{2}+13x+4\]

OpenStudy (tayfishb):

i know im supposed to factor now but i see no way of factoring

OpenStudy (tayfishb):

@Shadow_helper

OpenStudy (shadow_helper):

sorry dude ... i got no idea... try @YanaSidlinskiy

OpenStudy (tayfishb):

@TheSmartOne

OpenStudy (tayfishb):

@brainzonly

OpenStudy (sshayer):

\[f \prime(c)=\frac{ f(b)-f(a) }{ b-a }\]

OpenStudy (tayfishb):

@mathmale @Nnesha

OpenStudy (sshayer):

\[f \prime(x)=\frac{ (x+2)*1-x*1 }{ (x+2)^2 }=\frac{ 2 }{ (x+2)^2 }\] \[\frac{ 2 }{ (c+2)^2 }=\frac{ \frac{ 4 }{ (4+2) } -\frac{ 1 }{ (1+2) }}{ 4-1 }=\frac{ \frac{ 2 }{ 3 }-\frac{ 1 }{ 3 } }{ 3 }=\frac{ \frac{ 1 }{ 3 } }{ 3 }=\frac{ 1 }{ 9 }\] \[(c+2)^2=18,c+2=\pm \sqrt{18}=\pm3\sqrt{2},c=-2 \pm3\sqrt{2}\] now \[c=-2+3\sqrt{2}\in (1,4)\]

OpenStudy (sshayer):

\[c=-2+4.24=2.24 \in (1,4) \]

OpenStudy (tayfishb):

1) Where did c come from 2) why is it equal to 18 3)what did you do?

OpenStudy (sshayer):

read the statement of your question.

OpenStudy (sshayer):

c is value of x.

OpenStudy (tayfishb):

ok but why 18 and what came after that because i see what you did but i do not know why

OpenStudy (freckles):

You differentiated incorrectly at tayfish

OpenStudy (sshayer):

\[f \prime(c)=\frac{ 2 }{ (c+2)^2 }=\frac{ 1 }{ 9 }\] cross multiply \[(c+2)^2=9*2,c=2=\pm 3\sqrt{2}\]

OpenStudy (sshayer):

at c the tangent to curve is parallel to the chord joining the points on the curve where x=1 andx= 4

OpenStudy (sshayer):

correction above \[c=-2\pm 3\sqrt2 \]

OpenStudy (tayfishb):

ooohhhhhh oookkk thank you soo much!!!!

OpenStudy (sshayer):

yw